OpenStudy (vortish):

factor as a perfect square trinomial whenver possible 25^2-60x+36

OpenStudy (anonymous):

25^2-60x+36 = (5x)^2 - 2(5x)(6) + 6^2 Compare with \((a \pm b )^2 = a^2 \pm 2ab + b^2\) Factorize it.